Ryan is 4 years younger than Michael. The product of their ages is 1020 years.
What are both ages?

Answer by fcabanski(1391)

Check my profile and contact me for one on one online tutoring.


Michael is X years old.


Ryan is X-4 years old.


The product of their ages is 1020:


x(x-4)=1020 multiply both terms inside () by x:


x^2-4x=1020 (subtract 1020 from both sides):


x^2 - 4x -1020 =0 (factor with reverse foil)


(x + 30)(x - 34) (set each = 0 and solve):


x+30=0 ---> x=-30 age can't be negative, so this isn't the answer.


x-34=0 ---> x=34